100-100/100-100=2 0r 1/2 how?

Grade:upto college level

1 Answers

Sher Mohammad IIT Delhi
10 years ago
100-100=0=200-200=2(100-100)
100-100/100-100=2(100-100)/(100-100)=2
=(100-100)/2(100-100)=1/2
Sher Mohammad
B.Tech, IIT Delhi